OpenStudy (meilendurcer):

Is the simplified form of 2 √3 - 2√3 rational? I'm going with no.

ganeshie8 (ganeshie8):

wat do u get after simplification ?

OpenStudy (meilendurcer):

i would think 0

ganeshie8 (ganeshie8):

yup ! Isn't 0 rational ?

OpenStudy (meilendurcer):

oh thank you very much! Can i have you help with another?

ganeshie8 (ganeshie8):

shoot

ganeshie8 (ganeshie8):

btw, 0 IS a rational number !

OpenStudy (meilendurcer):

I have to simplify this:

ganeshie8 (ganeshie8):

simplify the last two terms first

ganeshie8 (ganeshie8):

\(\large \sqrt{45} = \sqrt{9 \times 5} = ?\)

OpenStudy (meilendurcer):

sqrt of 45 is 6.70?

ganeshie8 (ganeshie8):

dont use calculator :/

ganeshie8 (ganeshie8):

whats the square root of 9 ?

ganeshie8 (ganeshie8):

\(\large 3^2 = 9\), so \(\large \sqrt{9} = 3\)

ganeshie8 (ganeshie8):

use that to simplify^

ganeshie8 (ganeshie8):

\(\large \sqrt{9 \times 5} = 3\sqrt{5}\)

OpenStudy (meilendurcer):

i don't know how to do sqrt on calculator anyway. Sorry something came up. I'm back

ganeshie8 (ganeshie8):

\(\large 3\sqrt{5} - 2\sqrt{7} + \sqrt{45} - \sqrt{28}\) simplify last two terms : \(\large 3\sqrt{5} - 2\sqrt{7} + \sqrt{9 \times 5} - \sqrt{4 \times 7}\) \(\large 3\sqrt{5} - 2\sqrt{7} + 3\sqrt{5} - 2\sqrt{7}\)

ganeshie8 (ganeshie8):

next, combine like terms

ganeshie8 (ganeshie8):

\(\large 3\sqrt{5} - 2\sqrt{7} + \sqrt{45} - \sqrt{28}\) simplify last two terms : \(\large 3\sqrt{5} - 2\sqrt{7} + \sqrt{9 \times 5} - \sqrt{4 \times 7}\) \(\large 3\sqrt{5} - 2\sqrt{7} + 3\sqrt{5} - 2\sqrt{7}\) combine like terms : \(\large 6\sqrt{5} - 4\sqrt{7} \)

ganeshie8 (ganeshie8):

^thats the final simplified form

OpenStudy (meilendurcer):

can you be more specific? I can't do sqrt or cube roots at all

OpenStudy (meilendurcer):

that's the one i was going for too!! Thanks.